OpenStudy (anonymous):

HELP?!! Rationalize the denominator and simplify. a) 5/4- √3 b) 12/√7 + 2 c) 6/√2 - √3 d) 15√6/3√5

zepdrix (zepdrix):

Multiply the top and bottom by the `conjugate` of the denominator. \[\Large\bf\sf \frac{5}{4-\sqrt3}\left(\frac{4+\sqrt3}{4+\sqrt3}\right)\]Do you remember what a conjugate is? Does this step make sense?